### v1

# PsyMem: Fine-grained psychological alignment and Explicit Memory Control for Advanced Role-Playing LLMs

Abstract
Existing LLM-based role-playing methods often rely on superficial textual descriptions or simplistic metrics, inadequately modeling both intrinsic and extrinsic character dimensions. Additionally, they typically simulate character memory with implicit model knowledge or basic retrieval augment generation without explicit memory alignment, compromising memory consistency. The two issues weaken reliability of role-playing LLMs in several applications, such as trustworthy social simulation. To address these limitations, we propose PsyMem, a novel framework integrating fine-grained psychological attributes and explicit memory control for role-playing. PsyMem supplements textual descriptions with 26 psychological indicators to detailed model character. Additionally, PsyMem implements memory alignment training, explicitly trains the model to align character’s response with memory, thereby enabling dynamic memory-controlled responding during inference. By training Qwen2.5-7B-Instruct on our specially designed dataset (including 5,414 characters and 38,962 dialogues extracted from novels), the resulting model, termed as PsyMem-Qwen, outperforms baseline models in role-playing, achieving the best performance in human-likeness and character fidelity.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
1 Introduction
Recent advancements in large language models (LLMs) (Achiam et al., 2023; Anthropic, 2024a; Anil et al., 2023) have unlocked new possibilities for implementing role-playing systems through language-based behavior simulation (Shanahan et al., 2023). Role-playing systems demonstrate societal value across applications like interactive gaming (Wang et al., 2023a) and AI counseling (Zhang et al., 2024), particularly in enabling social simulation. Such simulation allows modeling of human interactions and societal dynamics (Park et al., 2023), offering insights into policy design and conflict mediation (Zhou* et al., 2024). To achieve trustworthy social simulation, LLM-based role-playing must preserve persistent traits, memories, and behavioral logic to ensure agents reliably mirror specified attributes in evolving social contexts.
Recent works in social or group behavior simulation (Kosinski, 2024; Park et al., 2023; Zhou* et al., 2024) often rely on general LLMs for role-playing, using context-based learning (Mann et al., 2020) and instruction follow-up (Ouyang et al., 2022). However, these approaches (Yu et al., 2022) commonly struggle to precisely control character attributes due to the unsatisfactory role-playing ability of general LLMs, especially for ordinary characters, as Fig.1 shows. Some recent works (Shao et al., 2023; Zhou et al., 2023; Tao et al., 2023) proposed the LLMs that are dedicated to role-playing to improve character attribute consistency. However, there are still the following two issues: 1) Oversimplified characterization and 2) Weak memory control.
The oversimplified characterization issue. From both neuroscientific and psychological perspectives, human attributes can be described through two broad dimensions: intrinsic attributes and extrinsic attributes (Snyder, 1983; Kahneman, 2011). Intrinsic attributes, such as personality traits and values (Ravlin and Meglino, 1987; Stumpf and Dunbar, 1991), play a significant role in decision-making, particularly in long-term decisions (Borghans et al., 2008; Roberts et al., 2008). On the other hand, extrinsic attributes (Kahneman, 2011), such as behavior, language, and body language, reflect an individual’s intuitive responses in immediate situations (Bargh et al., 1996).
However, existing role-playing LLMs typically rely on basic textual descriptions or narrow metrics to define the target role, which fails to capture the full complexity and depth of the character. For example, the work (Zhou et al., 2023) defines character attributes using textual descriptions based on seven dimensions (including identity, interests, experiences, and other relevant traits). Some recent research (Ran et al., 2024; Yu et al., 2024; Cui et al., 2023) has attempted to use MBTI for personality trait evaluation to improve character control. However, MBTI employs a binary approach to assess personality traits, which has low test-retest reliability (Stein and Swan, 2019; Hunsley et al., 2015) and cannot precisely quantify attributes (Nowack, 1996), such as how introverted a person is.
Figure 1: The performance comparison is conducted on two subsets: “Ordinary,” consisting of 20 randomly selected characters from our test set, and “Famous,” consisting of 20 well-known characters. The average attribute scores are calculated as the mean of the quantized scores across the 20 characters in each subset.
The weak memory control issue. Numerous neuroscientific studies have shown that(Stumpf and Dunbar, 1991; Ravlin and Meglino, 1987; Pickering and Garrod, 2004), during decision-making, the brain relies on not only personality traits (such as emotional responses and values) but also long-term memory, particularly in fast-paced decision-making contexts. However, existing role-playing models (Shao et al., 2023; Zhou et al., 2023) typically rely on the model’s internal knowledge to simulate character memory, which works for famous characters but often fails to provide memory for less familiar or unseen ones. Retrieval-Augmented Generation (RAG) (Ram et al., 2023) has been proposed as a promising solution (Zhong et al., 2024), but we observe that simply integrating RAG within Role-Play LLMs’ inference stage does not adequately address memory management in the context of role-play (see Table 2).
In this work, we propose a widely recognized modern psychological evaluation framework (Costa and McCrae, 2008; Schwartz, 1992; Thomas, 2008; Zimbardo and Ruch, 1975) and a graph-structured memory mechanism to construct the PsyMem framework, enabling LLMs to achieve exceptional role-playing abilities. Drawing from modern psychology and neuroscience, we enhance the description of character attributes by leveraging both latent psychological attributes and explicit behavioral patterns (Snyder, 1983; Kahneman, 2011), with personality and values corresponding to latent psychological attributes, and behavioral decision-making, social interaction, and leadership representing explicit behavioral patterns. Overall, we describe characters using 26 quantitative indicators and a small amount of textual descriptions. We extract real data from novels and construct a publicly available role-playing dataset. As shown in Table 1, this dataset contains 5,414 established characters, 38,962 conversations, and 536,636 utterances, much larger than existing role-playing datasets.
For character memory, inspired by the hippocampus and cognitive mapping mechanisms (Eichenbaum, 2015; Hartley et al., 2014; Tacikowski et al., 2024; Tavares et al., 2015), we transform characters, events, and relationships into a knowledge graph. Unlike previous role playing work (Edge et al., 2024) directly using implicit model knowledge or integrating role memory in inference, we incorporate memory information into the training process, forcing the model to learn how to respond given retrieved memory. This approach effectively enhances the model’s memory consistency, ensuring that role-playing model generates responses based on not only character attributes but also memory.
In general, our five main contributions are as follows:
  • We propose PsyMem, a novel LLM-based role-playing framework that establishes precise control over character behaviors through the synergistic integration of fine-grained psychological profiling (e.g., personality, value) and dynamic memory schemata.
  • Beyond textual descriptions, we present a comprehensive set of attributes that bridges intrinsic (e.g., values, emotions) and extrinsic (e.g., behaviors, social interactions) character dimensions, operationalized via 26 measurable indicators. This attribute set captures both the internal motivations and external behavioral manifestations of characters.
  • We introduce a memory alignment training paradigm that establishes explicit cognitive anchoring between role-playing responses and character memory. This methodology enhances role-playing through a dual-control mechanism, enabling precise regulation of generative outputs via both trait-based attributes (e.g., personality dimensions) and episodic memory constraints.
  • We introduce the use of role-based general supervised fine-tuning data (transforming general SFT data into role-playing style SFT data), demonstrating that incorporating such data effectively enhances the performance of Role-Playing LLMs.
  • Extensive experiments demonstrate that the proposed PsyMem framework significantly improves the role-playing ability of LLMs. For instance, by training Qwen2.5-7B-Instruct on our dataset (introduced in Section 3) specially designed for PsyMem, the resulting model, termed as PsyMem-Qwen, outperforms baseline models in role-playing, achieving the best performance in human-likeness and character fidelity.
2 Related Work
2.1 LLMs-based Role-Playing
In existing role-playing works, the main approaches are Nonparametric Prompting and Parametric Training (Chen et al., 2024). In Nonparametric Prompting, Yu et al. (2022) use carefully designed character descriptions to control general LLMs for role-playing without fine-tuning. However, they face significant challenges in accurately reflecting the intrinsic relationships between character profiles and dialogue content.
In Parametric Training, Chen et al. (2023) focus on mimicking characters from the Harry Potter series, while Wang et al. (2023b) introduced the first role-playing dataset with 100 characters generated using GPT-3.5 prompts. Subsequent works (Li et al., 2023; Zhou et al., 2023; Shao et al., 2023) have expanded and enriched these datasets by improving the number of characters and the diversity of dialogue scenarios, all leveraging various GPT models. Meanwhile, Lu et al. (2024) adopted a self-alignment approach to dataset construction, moving away from the previous method of cheaply imitating the role-playing capabilities of GPT models. Zhou et al. (2023) leveraged GPT-4 to extract character dialogues from a diverse collection of Chinese novels and scripts. More recently, Yu et al. (2024) focused on addressing the bias between predefined characters and specific scenario dialogues, and constructed a dataset of 311 characters.
Our work belongs to parametric Training. We argue that a character’s behavior is jointly influenced by their attributes, memory, and the scenario. Drawing from modern psychology, we segment character attributes using contemporary psychological frameworks and construct a dataset containing 38,962 conversations (536,636 utterances) from real novels, leveraging both character attributes and memory to enhance the model’s role-playing ability.
2.2 Memory Mechanism in Role-Playing.
Existing LLMs-based Role-Playing commonly relies on the internal knowledge within the model weights to implicitly mimic character memory (Zhou et al., 2023; Lu et al., 2024; Li et al., 2023). For example, Shao et al. (2023) and others push the boundaries of knowledge to protect character memory, while Lu et al. (2024) and others use a self-alignment approach to distinguish the different memory backgrounds of various characters. Relying on the internal knowledge within the model is effective for famous characters but fails for infamous and unseen characters. Some works apply Retrieval-Augmented Generation (RAG) (Ram et al., 2023) which dynamically retrieving external contexts to explicitly mimic character’s memory. For example, Zhong et al. (2024) adopts a memory update mechanism inspired by Ebbinghaus’s forgetting curve theory, emphasizing the recall of relevant memories to understand the user’s personality. However, as shown in Table 2, existing role-playing models tend to perform better when simulating famous characters, while their performance significantly drops for ordinary characters that rely on external memory.
Unlike previous works, we integrate explicit memory control into the training framework of role-playing models. This approach encourages the model to strictly rely on actual retrieved memories during dialogue generation in inference.
2.3 Role-playing Evaluation.
The current evaluation process consists of dialogue generation and dialogue assessment. Due to its efficiency and completeness, automated dialogue generation is widely adopted. Evaluation methods are generally categorized into three types: Metric-based Evaluation(Li et al., 2023; Wang et al., 2023b), Human Evaluation(Zhou et al., 2023), and “LLMs as Judges”(Shao et al., 2023; Lu et al., 2024). Metric-based evaluation primarily measures the model’s alignment with standard responses, such as original text from novels or manually annotated content. While human evaluation is more precise, its high cost and lack of reproducibility limit its scalability. In contrast, “LLMs as Judges”, with its efficiency, low cost, and strong scalability, is increasingly becoming the preferred choice.
3 Dataset Architecture
An increasing number of role-playing datasets (Ran et al., 2024; Cui et al., 2023) have begun incorporating modern psychological indicators into their characterization frameworks. However, previous studies typically introduced only a limited selection of psychological attributes as supplementary descriptors, without comprehensively integrating psychological systems into character portrayals (Yu et al., 2024; Wang et al., 2024). Moreover, existing datasets predominantly focus on character attributes, overlooking the significant influence memory has on characters’ behaviors, thought processes, and speech styles(Stumpf and Dunbar, 1991; Ravlin and Meglino, 1987; Pickering and Garrod, 2004). This limitation becomes particularly evident when role-playing characters possess limited background knowledge during the pretraining stage (see Figure 1).
To address these shortcomings, we design a novel dataset with the architecture formulated as
[…]
where $N$ is the number of data sample. Each sample is a dialogue turn where $R_{i}$ is the role for response; $P_{i}$ is the profile of $R_{i}$, featuring 26 quantitative psychological dimensions; $C_{i}$ is the previous dialogue turns before the current dialogue. $x_{i}$ is the query to $R_{i}$ and $y_{i}$ is the corresponding response. $M_{i}$ is the retrieved memory response of role $R_{i}$ with respect to query $x_{i}$ and the previous dialogue $C_{i}$.
In this work, the dataset $D_{\mathrm{RP}}$ contains 38,962 samples, extracted from 539 novels characterized by a wide range of genres, rich content, and narrative perspectives of third person.(The genre distribution is presented in Figure 2.)
Dataset Character Conversation Profile Style
Num. SA Conv. Avg. Turns Auth. PsyFrm QI Text Desc. Expl. Mem.
HPD 113 1,191 13.2
CharacterGLM 250 1,034 15.8
RoleLLM 100 140,726 2
CharacterLLM 9 14,300 13.2
ChatHaruhi 32 54,726 >2
DITTO 4002 7,186 5.1
Beyond Dialogue 311 3,552 6.5
CoSER 17,966 29,798 13.2
MMRole 85 14,346 4.2
CharacterBench 3956 13,162 11.3
CharacterEval 77 1,785 9.3
Ours 5,414 38,962 13.8
Table 1: Dataset Statistics. Comparison between our dataset and existing open-source role-play datasets. Columns are categorized into: Character (Num.: number of character roles; SA: scene alignment), Conversation (Conv.: number of dialogues; Avg. Turns: average dialogue length; Auth.: whether dialogues are fully sourced from real, non-generated conversations), and Profile Style (PsyFrm: adherence to a psychological framework; QI: presence of quantitative indicators; Text Desc.: availability of textual character descriptions; Expl. Mem.: support for explicit memory retrieval during interaction).
Figure 2: The genre distribution in the dataset.
Figure 3: (a): The fine-grained character attributes designed in this work. (b): The two stages training of role-playing LLM. First, the model learns basic role-playing without character memory. In the second stage, we dynamically retrieves memory relevant to the current query and dialogue context from graph-structured character memory. We then integrate character profile, retrieved memories, dialogue history, and current input to enhance role-playing precision, training the model to align responses with both fine-grained character profile and contextual memory. (c):We assess the role-playing LLM by first generating a multi-turn dialogue (up to 15 turns) between two designated roles, followed by three rounds of scoring with GPT-4o based on a quantitative rubric, and report the mean score.
3.1 Quantifiable Character Attributes
We systematically construct quantifiable character attribute dimensions aligned with contemporary psychological frameworks by categorizing character traits into latent psychological attributes and explicit behavioral patterns (Kahneman, 2011; Snyder, 1983), as illustrated in Figure 3a. These two dimensions distinctly influence individual behavior, communication styles, decision-making processes, and interpersonal interactions. Besides, a small amount of text describing the social attributes of the characters is also included as a supplement.
Latent psychological attributes reflect internal, often subconscious aspects of character, guiding motivations, preferences, and emotional responses. To describe these latent psychological attributes comprehensively, we utilize the following psychological frameworks:
1) Personality: Evaluated using the Big Five Personality Model (Costa and McCrae, 2008), encompassing stable emotional, cognitive, and behavioral tendencies. Traits such as openness, conscientiousness, extraversion, agreeableness, and neuroticism directly influence communication styles, social interactions, and general life choices.
2) Values: Assessed through Schwartz’s Theory of Basic Values (Schwartz, 1992), which classifies core beliefs into ten universally recognized types. These values deeply affect decision-making, priorities, ethical judgments, and interpersonal alignment, enriching the psychological depth and motivational dimensions of individuals.
Explicit behavioral patterns manifest externally in observable actions, interactions, and communication styles. We capture these explicit patterns using:
1) Social & Leadership Styles: Measured through six dimensions based on Zimbardo’s psychological principles (Zimbardo and Ruch, 1975), significantly influencing observable interpersonal behaviors, communication effectiveness, leadership approaches, and social dynamics in groups.
2) Behavioral Decision: Analyzed using the Thomas-Kilmann Conflict Mode Instrument (TKI) (Thomas, 2008), emphasizing decision-making processes when characters face conflicts or challenges. This explicitly captures how emotional, social, and heuristic influences shape specific decisions, providing precise and actionable insights into behavioral responses.
3.2 Character Memory Integration
Neuroscientific research underscores the significant role memory plays in shaping human behavior, communication, and decision-making processes (Stumpf and Dunbar, 1991; Ravlin and Meglino, 1987; Pickering and Garrod, 2004). Memories influence not only how individuals perceive and interpret events but also guide their interactions and responses within social contexts, highlighting the importance of accurately capturing and utilizing memory in character portrayal.
Therefore, we introduce character memory data $M$ to the role playing dataset to enable the memory simulation capability of role playing model. Specifically, we first build a knowledge graph based on the novel via GraphRAG (Edge et al., 2024), and then retrieve relevant memory for the current dialogue via entity mentions, relational triples, and raw dialogue chunks associated with specific scenes. Memory alignment is evaluated on two key dimensions: correctness and rationality. Correctness assesses whether the model accurately utilizes stored character information in response to queries, while rationality examines logical coherence and contextual appropriateness. Additionally, simple summaries or vague memory references are identified as non-rational uses of memory. To further enhance robustness, irrelevant memory items are deliberately introduced as noise, promoting stable, realistic, and consistent character portrayals reflective of human memory dynamics.
In addition, we define character-specific knowledge boundaries to ensure consistency—for example, Quasimodo from The Hunchback of Notre-Dame would not be aware of torch. Based on the absence of relevant retrievable memories, we used GPT-4o to generate 800 refusal-style QA pairs for 100 characters, grounded in real-world constraints and narrative context. These were integrated into novel-based dialogues to maintain coherence and character fidelity.
3.3 Role-based SFT Data
We propose a Role-based general supervised fine-tuning (SFT) data construction approach. Dialogue samples from the Pure-Dove (PD) dataset (Daniele and Suphavadeeprasit, 2023) are grouped randomly, and LLMs infer each speaker’s latent psychological attributes and explicit behavioral patterns. Scene-specific information is dynamically generated using fine-tuned LLMs based on dialogue contexts, significantly improving character consistency and alignment during interactions, as demonstrated by our experimental evaluations.
Similar to Equation 1, we construct a dataset defined as:
[…]
Where $R_{i}$ denotes the role responsible for the response, and $P_{i}$ is the profile of $R_{i}$ defined by 26 quantitative psychological dimensions. $C_{i}$ represents the preceding dialogue context, $x_{i}$ is the current query directed to $R_{i}$, and $y_{i}$ is the corresponding response. These $(x_{i},C_{i},y_{i})$ triples are drawn from general supervised fine-tuning datasets. In addition, $M_{i}$ refers to the retrieved memory response relevant to role $R_{i}$ with respect to $x_{i}$ and $C_{i}$.
4 Role Playing Training and Evaluation
4.1 Training
In this work, we transfer general LLMs to role-playing models via supervised fine-tuning. Role-playing presents a unique challenge for language models: maintaining both general capabilities and character-specific attributes simultaneously. This challenge exemplifies the fundamental tension in transfer learning between preserving general knowledge and specializing for domain-specific tasks. Traditional fine-tuning approaches (Kotha et al., 2024) often suffer from catastrophic forgetting, where optimization for character fidelity compromises the model’s general capabilities. Inspired by the Dual-stage Mixed Fine-tuning (DMT) (Dong et al., 2023) method, we introduce a two-stage training strategy for role-playing, effectively balance character fidelity and the preservation of general language understanding capabilities, as illustrated in Figure 3b.
Specifically, we introduce a two-stage progressive training framework to systematically develop LLMs into specialized Role-Playing Models: Stage 1 establishes core role-playing capabilities through fine-tuning on data emphasizing fundamental role characteristics. Stage 2 enhances specialization via memory-augmented fine-tuning, incorporating long-term memory contexts and general role-based information. This approach ensures a balanced development of character-specific attributes while simultaneously preserving the model’s general capabilities.
Stage 1: Foundational Role-Playing Capacity Development.
In this stage, we train LLMs to learn foundational role-playing ability based only on quantifiable attributes and basic information. We fine-tune the model using a subset $D_{\mathrm{RP1}}$ sampled from the overall role-playing dataset to establish its fundamental role-playing abilities. The objective function is formally defined as:
[…]
where ()($\cdot$)( ⋅ ) in $D_{\mathrm{RP}}$ represents the tuple $(R,P,C,x,y)$, $D_{\mathrm{RP1}}$ is a subset sampled from the overall role-playing dataset $D_{\mathrm{RP}}$, excluding the retrieved memory $M$.
Stage 2: Integrated Memory-Augmented Specialization. To achieve integrated memory-augmented specialization, we further specialize the model by fine-tuning it on a distinct subset from the role-playing dataset $D_{\mathrm{RP}}$, augmented with long-term memory contexts retrieved via graph-based methods, and combined with general role-based SFT data $D_{\mathrm{RB}}$. This stage employs a weighted combination of two loss terms, explicitly designed to mitigate forgetting of general capabilities while simultaneously enhancing the model’s ability to effectively leverage long-term memory contexts:
[…]
where $\alpha=20$.
The two loss terms are individually defined as follows:
[…]
where ()($\cdot$)( ⋅ ) in $D_{\mathrm{RP2}}$ represents the tuple $(R,P,M,C,x,y)$ and in $D_{\mathrm{RB}}$ represents $(R,P,C,x,y)$. $M$ denotes the retrieved memory response for role $R$ regarding the query $x$ and prior dialogue context $C$, obtained via graph-based retrieval methods and subsequently filtered by LLMs.
Model Character Fidelity Character-independent
Per. Val. SL BD Mem. Avg. H-like Cons. Coh.
General LLMs
GPT-4o 79.08 80.11 77.43 73.29 94.40 80.86 66.60 89.28 99.20
GPT-3.5-Turbo 79.08 79.42 74.15 72.11 84.60 77.87 34.40 88.60 96.20
Claude-3.5-sonnet 80.34 80.29 78.51 74.57 91.80 81.10 75.00 89.50 99.20
Qwen-Max 78.67 80.42 76.54 72.37 90.60 79.72 57.60 89.48 97.60
Yi-Large 80.20 80.09 78.27 76.82 92.00 81.48 83.80 88.18 98.40
Deepseek-R1 80.98 81.38 77.88 77.12 95.00 82.47 77.20 87.57 96.20
Deepseek-V3 79.82 80.33 79.16 75.08 93.40 81.56 76.40 89.04 99.00
Llama3.1-8B-instruct 79.17 80.14 76.76 76.01 83.20 79.06 64.20 88.74 99.00
Qwen2.5-7B-Instruct 77.44 80.02 76.15 76.51 85.60 79.14 64.40 89.58 97.40
Role-play LLMs
CharacterGLM-6B 80.66 80.67 79.35 71.20 40.20 70.42 52.80 89.61 73.00
Baichuan-NPC-Turbo 76.53 79.05 74.29 72.27 87.20 77.87 61.60 90.36 96.40
Hunyuan-Role 79.31 79.71 78.24 78.38 82.00 79.53 87.40 89.16 91.20
Higgs-Llama3-70B 80.02 80.61 75.77 74.08 93.20 80.74 60.40 88.40 99.60
CoSER-70B 81.21 81.90 78.83 76.35 93.00 82.28 81.80 89.44 97.20
PsyMem-LLama (ours) 80.47+1.30 81.93+1.79 78.57+1.81 74.47-1.54 90.20+7.00 81.13+2.07 85.20+21.00 89.93+1.19 95.60-3.40
PsyMem-Qwen (ours) 80.40+2.96 81.74+1.72 80.80+4.65 78.48+1.97 91.80+6.20 82.64+3.50 87.60+23.20 89.64+0.06 96.20-1.20
Table 2: Model evaluation across multiple criteria. Bold indicates best performance. Superscripts show change from base model. Abbreviations: Per. = Personality, Val. = Values, SL = Social & Leadership, BD = Behavioral Decision, Mem. = Memory, H-like = Human-likeness, Cons. = Consistency, Coh. = Coherence
4.2 Evaluation
We evaluate role-playing capabilities of LLMs by adopting the widely used “LLMs as Judges” approach (Shao et al., 2023; Lu et al., 2024), as illustrated in Figure 3c. The evaluation process is structured into two main categories: Character-independent Capabilities and Character Fidelity. Additionally, in the Tabel 4, we further discuss the discrepancies between LLM-based judgments and human evaluations across 26 quantitative indicators to validate the effectiveness of the “LLMs as Judges” approach.
4.2.1 Character-independent Capabilities
Following previous research(Chen et al., 2024; Yu et al., 2024), we measure Character-independent Capabilities using three metrics: Human-likeness, Consistency and Coherence. In the Human-likeness task, we use GPT-4o with few-shot examples to determine whether a dialogue was generated by a human, ultimately computing the accuracy score. For consistency, we evaluate the coherence between different parts of the conversation to assess whether the role-playing remains consistent throughout long dialogues. Additionally, for coherence, we assess how logically connected and contextually appropriate the responses are across the entire dialogue, ensuring the conversation flows naturally.
4.2.2 Character Fidelity
Character Fidelity assesses how accurately an LLM portrays specific role-playing characters. As described in Section 3.1, we employ modern psychological quantification frameworks defined in our dataset design, covering four primary dimensions: Personality (Big Five Personality Model), Values (Schwartz’s Theory of Basic Values), Social and Leadership Styles (Zimbardo’s psychological principles), and Behavioral Decision-making (Thomas-Kilmann Conflict Mode Instrument, TKI). These dimensions are quantified into numeric scales using standardized psychological evaluation methods. Additionally, we separately evaluate the model’s memory, focusing on the correctness and rationality of memory alignment. The Memory dimension is measured using accuracy scores derived from a true/false evaluation format.
To calculate similarity between role-playing characters, we employ Normalized Euclidean Distance (NED), computed as follows:
[…]
where $\Delta X=X_{\max}-X_{\min}$, $n$ is the number of subdimensions, $X_{\mathrm{ref}}$ is the standardized reference value, and $X_{i}$ represents the measurement for the $i$-th subdimension, with possible values from $Z=\{1,2,3,4,5\}$. This formulation is robust, as it normalizes the mean squared deviation across different scales, yielding a more interpretable similarity metric.
5 Experiments
In this section, we explain the dataset, baselines, and metrics design in the experimental setup, and discuss the main results. Additional ablation experiments are discussed in Section 5.3.
5.1 Experimental Setup
Evaluation Dataset.
To mitigate potential knowledge bias introduced during the pretraining phase, we construct an evaluation dataset for our main experiments by selecting characters and scenes from novels published after June 2024. This design choice aims to ensure that models have had minimal or no exposure to the selected texts during pretraining, thereby reducing reliance on memorized character-specific knowledge. We first score 100 contemporary novels using large language models (LLMs) and then select 50 representative characters spanning diverse genres. For each character, we collect ten scenes directly sourced from the original text.
Baselines.
In our benchmark evaluation, we tested both open-source and proprietary advanced chatbots, each representing the state-of-the-art within their respective frameworks. For general-purpose models, we compared the proposed PsyMem with several prominent baselines, including GPT-4o(Hurst et al., 2024), GPT-3.5-turbo (OpenAI, 2023),Claude-3.5-sonnet(Anthropic, 2024b), Qwen-Max (Team, 2024), Yi-Large (Young et al., 2024), Deepseek-R1(Guo et al., 2025), DeepSeek-V3 (Liu et al., 2024), Llama3.1-8B-instruct(Grattafiori et al., 2024) and Qwen2.5-7B-Instruct (Yang et al., 2024). These models are considered some of the best in the field, offering a broad range of capabilities across various domains. In the category of Role-Playing-focused baselines, we made comparisons with Hunyuan-Role(Tencent Cloud, 2023), Baichuan-NPC-Turbo(Baichuan AI, 2023), and CharacterGLM-6B (Zhou et al., 2023),Higgs-Llama3-70B(AI, 2024),CoSER-70B(Wang et al., 2025), which are specifically designed to handle character-driven interactions and story-based tasks.
Implementations.
With the proposed PsyMem framework, we train our models based on Llama3.1-8B-Instruct and Qwen2.5-7B-Instruct, and denote the trained models as PsyMem-LLama and PsyMem-Qwen, respectively. All models were trained on four A800 80G GPUs for 3 epochs with a global batch size of 128 and a sequence length of 8,192. LoRA was used with a per-device batch size of 2 and 16 gradient accumulation steps. We adopted a cosine learning rate schedule with a warmup ratio of 0.1, and set the learning rate to 1e-4.
Metrics
We utilize GPT-4o to evaluate 50 established characters across 500 scenarios, each scenario comprising 15 dialogue turns. Evaluations are conducted based on the dimensions described in Section 3, with mean scores calculated accordingly.
5.2 Main Results
We present the main results in Table 2. We extracted 500 real novel scenes to evaluate all models, assessing them from two perspectives: Character Fidelity and Character-independent Capabilities. Notably, in terms of Memory alignment, Proprietary models demonstrate a more reasonable approach—not merely summarizing but leveraging memory to portray characters more authentically. Although Higgs-LLaMA3-70B and CoSER-70B exhibit relatively stronger performance in memory handling compared to other role-play-specialized models, there remains a noticeable gap between them and top-performing general-purpose models. Within the General Baselines, we observe that Proprietary models still significantly outperform open-source models. GPT-4o and DeepSeek-V3, as outstanding proprietary models, exhibit impressive performance in Character-independent Capabilities. Additionally, Baichuan-NPC-Turbo, as a role-play expertise baseline model, surpasses general baselines in self-consistency, indicating that this metric is particularly important for role-playing agents. Furthermore, it is worth mentioning that DeepSeek-R1 and CoSER-70B excel in Character Fidelity.
Empirical results under the PsyMem framework reveal that, relative to their original counterparts, PsyMem-Qwen and PsyMem-LLaMA achieve character fidelity improvements of 2.07% and 3.50%, respectively. In addition, PsyMem-Qwen demonstrates strong performance in the Social & Leadership and Behavioral Decision dimensions, achieving scores of 80.80% and 78.48%, respectively. PsyMem-LLaMA, on the other hand, obtains a score of 81.93% in the Value dimension, surpassing all other evaluated models in this category. Moreover, PsyMem-Qwen, despite having only 7B parameters, surpassed all measured baselines in Character Fidelity. Additionally, it is worth noting that both models exhibited remarkable improvements in Human-likeness. Overall, the PsyMem framework has demonstrated strong effectiveness in our benchmark tests, enabling precise character control through quantitative character attribute metrics and memory mechanisms.
Model Character Fidelity
Per. Val. SL* BD* Mem. Avg.
Llama3.1-8B-Instruct
+ PsyMem (w/o Mem.) 80.11 81.90 79.23 76.36 44.20 72.36
+ PsyMem 80.47 81.93 78.57 74.47 90.20 81.13
Qwen2.5-7B-Instruct
+ PsyMem (w/o Mem.) 81.02 81.33 82.10 79.02 53.00 75.29
+ PsyMem 80.40 81.74 80.80 78.48 91.80 82.64
Table 3: Ablation study on memory integration in the PsyMem framework. Per.: Personality, Val.: Values, SL: Social & Leadership, BD: Behavioral Decision, Mem.: Memory.
Metrics Per. Val. SL* BD*
Cosine 0.98 0.96 0.97 0.97
Table 4: Cosine similarity between human and GPT-4o scores. Per.: Personality, Val.: Values, SL: Social & Leadership, BD: Behavioral Decision.
5.3 Ablation Study
Model Character Human Consis- Coher-
Fidelity -likeness tency ence
Llama3.1-8B-Instruct 79.06 64.20 88.74 99.00
     + PsyMem (Original PD data) 80.51 70.80 89.49 82.20
     + PsyMem (Role-based PD data) 81.13 85.20 89.93 95.60
Qwen2.5-7B-Instruct 79.14 64.40 89.58 97.40
     + PsyMem (Original PD data) 81.63 86.60 89.45 93.40
     + PsyMem (Role-based PD data) 82.64 87.60 89.64 96.20
Table 5: Ablation study on Role-based (RB) Pure Dove (PD) dataset in the PsyMem framework.
Figure 4: The impact of data size on model performance using the Pure Dove (PD) dataset.
We further analyzed the effectiveness of memory injection and the incorporation of role-based general SFT data in training.
Memory alignment training. To evaluate the impact of memory alignment training on character fidelity, we constructed a memory testing dataset using GPT-4o, designing 1,500 questions across 50 roles in 500 scenarios based on detailed character profiles. We present our results in Table 3, comparing the performance of PsyMem-LLama and PsyMem-Qwen, with and without memory alignment training. Our findings indicate that incorporating memory data in the second training stage significantly enhances the model’s ability to responding based on retrieved memory, rather than merely summarizing relevant content. This improves both character attribute consistency and long-term memory retention. However, we also observe that memory alignment training has a slight negative impact on character attribute alignment. We hypothesize that this may be due to the presence of a small number of role-specific dialogue examples in the memory data, which could slightly weaken the model’s ability to strictly align with predefined character attributes.
Role-Based General SFT Data We investigate the impact of Role-based General SFT Data on Character-Independent Capabilities and Character Fidelity in Table 5 and Figure 4. Table 5 presents the ablation experiment on the use of Role-based General SFT Data, showing that PsyMem-LLama and PsyMem-Qwen helps mitigate the decline in coherence capabilities caused by catastrophic forgetting of general capabilities. Additionally, it leads to improvements in Character Fidelity, Human-likeness, and Consistency.
In Figure 4, we examine the impact of General SFT Data and Role-based General SFT Data on Character Fidelity for PsyMem-Qwen. The results validate the effectiveness of Role-based General SFT Data. Specifically, we observe that the use of General SFT Data negatively impacts the model’s character-specific abilities, whereas training with Role-based General SFT Data achieves optimal performance with 1,500 samples, after which performance begins to decline. Despite this, it consistently outperforms General SFT Data. We hypothesize that the decline in performance with 3,000 samples of Role-based General SFT Data occurs because the dataset’s distribution becomes overly dominated by GPT-generated content, causing an imbalance in role-specific attributes. A balanced amount of Role-based General SFT Data thus helps improve character-playing abilities without leading to such issues.
6 Conclusion
In this work, we propose a novel framework PsyMem for controlling character role-playing in large language models (LLMs) by integrating two key aspects: character memory and character attributes, both of which align more closely with modern psychological theories. We validate the proposed PsyMem using a large scale specially designed role-playing dataset based on real novels, addressing the gap between generic and well-known characters in role-playing tasks. Our experiments show that PsyMem-Qwen, trained with our alignment across five dimensions, outperforms all baselines mentioned in this paper, demonstrating strong performance in Character-Independent Capabilities as well.
Furthermore, we explore the benefits of incorporating Role-based General SFT Data for role-playing models. Through this work, we aim to encourage future research on role-playing models that consider psychological frameworks for a more nuanced understanding of character behavior, ultimately advancing the development of more realistic and consistent role-playing LLMs.